Find the voltage across a resistor if the current flowing through it is 5A. And also determine the power and energy consumed by the resistor if the current is allowed to pass through the resistor for 20 min. 


1
Expert's answer
2021-04-19T18:03:07-0400

To be given in question

Current (I)= 5Amp

Time(t)=20 minutes="\frac{20}{60}=\frac{1}{3} hour"

To be asked in question

Power (P)=?

Energy (E)=?

We know that

Power

"P=I^2R\rightarrow(1)"

Put value

"P=5^2\times R"

"P=\frac{25R}{1000}(kilowatt)"

"P=0.025R (kilowatt) \rightarrow(2)"

Energy

"E=P\times\frac {1}{3} \rightarrow(3)"


Equation (2) ,(3)we can written as

"E=0.025\times\frac {1}{3}\times R"

Solve it

"E=8.333R\times 10^{-3}(kilowatthour)"